1A of current produces a deflection of 30o
in a tangent galvanometer then its reduction
factor is
a) 1.414 b) 0.707 c) 1.732 d) 3
1

Grade:

2 Answers

Saurabh Koranglekar
3 years ago
Dear student

A reduction factor of a tangent galvanometer is equal tocurrent, required to be passed through the galvanometer to produce a deflection of 45∘. The reduction factor is measured in ampere

Regards
Vikas TU
3 years ago
Dear student 
i∝tanφ⇒i1/i2=tanφ1/tanφ2
⇒0.1/i2=tan30∘/tan60∘=1/3⇒i2=0.3A
Good Luck …...............................................................